# Search-Based Software Engineering in the Landscape of AI Foundation Models

### v1

@@ 2.3 Solution Encoding for SBSE Problems
A fundamental step in SBSE is solution encoding for a particular SE problem. This varies from one problem nature to another and requires problem-domain knowledge. FMs with diverse capabilities can support solution encoding for different SE lifecycle phases. For instance, to verify software models with SBSE during the design phase, MMs with both textual and visual capabilities can facilitate solution encoding by analyzing software requirements and design artifacts, often created using graphical modeling languages such as Unified Modeling Language (UML).
2.4 FMs for SBSE Implementation Generation
@@
One advanced application of FMs is generating either partial or complete implementations of SBSE problems. In a partial implementation, FMs generate code to solve an SBSE problem, which the user then completes manually. Alternatively, FMs can create complete implementations of SBSE problems that can be further refined. This capability extends across key SBSE components, including solution encoding, operator selection, fitness function design, and search algorithm selection, enabling more efficient and automated problem-solving.
2.5 FMs for Testing, Debugging, and Repair of SBSE Implementation
@@
3 Foundation Models for SE Artifacts Generated with Search-based Software Engineering
@@
In recent years, there has been growing research interest in various SE areas in using LLMs for SBSE. One notable area is software genetic improvement (GI). In this context, Kang and Yoo [20] presented the initial concept of generating mutants from LLMs to optimize the process of GI. Similar to this work, Brownlee et al. [5] employed LLMs as mutation operators to increase diversity in the search process for GI. Software program refactoring represents another emerging direction where FMs are applied in SBSE. In this regard, Choi et al. [10] utilized LLMs to guide a search-based approach, enhancing refactoring efficiency while reducing complexity. Another key area of interest is software testing, where LLMs are being explored for search-based software testing (SBST). In this case, Turhan [40] introduced evoLLve’M, an approach that utilizes LLMs to enhance JUnit test assertions generated by EvoSuite’s evolutionary process. Similar to this work, Biagiola et al. [4] employed LLMs to improve the readability of unit tests generated by EvoSuite through search techniques.
@@
In summary, the literature review indicates that (i) LLMs, among various FMs, are increasingly used for SBSE, and (ii) numerous SE phases, including requirements engineering, software design, and debugging, remain largely unexplored. Despite SBSE’s long-standing role in solving SE problems, several challenges persist, which we anticipate can be addressed through FMs in the future. Below, we highlight key opportunities and research directions.
3.1 FMs for SE Artifact Interpretation
@@
An SE problem in SBSE is formulated using different solution encoding schemes, such as Binary or Tree-based encoding, depending on the problem’s nature. The final output is generated as an encoded solution, often requiring manual effort for interpretation and transformation into a specific SE artifact. This effort varies based on the SE lifecycle phase. For instance, in the case of requirements prioritization for the next release problem (NRP), the SBSE-generated prioritization sequence needs to be interpreted in textural requirements and utilized for NRP. Similarly, software models generated with SBSE need to be converted to graphical models. In this context, FMs can play a significant role in two directions. First, LLMs with text comprehension and analytical capabilities can be used for SBSE-generated SE artifacts that are required to be interpreted and represented in a textual format. Second, VLMs and MMs can be employed to analyze and transform SBSE-generated SE artifacts into graphical representations.
3.2 FMs for SE Artifact Refinement
@@
For some SE problems, SBSE-generated solutions do not need to be transformed. For example, test data generation for primitive types and program improvement with generic programming. Such scenarios require domain experts to verify the generated solution manually. In the case of test data generation, test effectiveness using metrics like coverage needs to be analyzed after running tests on the system under test. Similarly, programs generated with generic programming either need to be compiled and tested, or they need to be manually examined using code inspection and walkthrough techniques. FMs have significant potential in these cases. For the test data generated using SBSE, LLMs can conduct a static analysis utilizing the source code and test data to provide insights on test effectiveness, such as untested paths or incomplete coverage scenarios. In developing LLM-based static analysis techniques, we foresee the potential for devising novel test effectiveness metrics that complement these techniques. Furthermore, LLM-based static analysis (e.g., type checking and data flow analysis) can be further extended to verify programs generated with SBSE statically.
3.3 FMs for Pareto Analysis
@@
In the case of multi- and many-objective SE problems, SBSE generates non-dominated or Pareto-optimal solutions in the objective space. A common practice is to analyze the resultant solutions and select the best one meeting a particular set of requirements. This process needs to be repeated whenever requirements change. In this regard, LLMs can support domain experts in understanding the trade-offs between different objectives and selecting the best solution for a given scenario. Moreover, for non-expert users, VLMs or MMs can construct a visual representation of the Pareto front, illustrating the set of all non-dominated solutions and offering suggestions to assist in selecting the most suitable option.
4 Search-based Software Engineering for SE Artifacts Generated with Foundation Models
FMs are increasingly being integrated into the entire SE lifecycle [23]. This results in FMs customized for specialized SE activities. A prominent example is LLMs tailored for different coding tasks, such as CodeX [9] for code generation and CodeGen [9] for program synthesis. In addition to code generation, these LLMs are employed to generate test cases [13]. For debugging, LLMs have been utilized for fault localization [43] and software bug reproduction from bug reports using prompting techniques [21]. Another line of research is focused on applying SBSE techniques to optimize FMs for specific SE problems. A noteworthy area is software effort estimation, where Tawosi et al. [39] introduced the idea of applying search techniques to optimize few-shot methods for fine-tuning LLMs to enhance their performance in estimating software effort. Another area of research focuses on enhancing image quality through SBSE. In this regard, Berger et al. [3] presented StableYolo, which employs a multi-objective search to optimize LLM prompts and parameters for generating high-quality and realistic images using tools such as YOLO and Stable Diffusion. Later, Gong et al. [16] introduced GreenStableYolo, which applies a multi-objective search for efficient LLM inference while maintaining high-quality image generation.
@@ 4 Search-based Software Engineering for SE Artifacts Generated with Foundation Models
A key limitation noted in the literature is the lack of sufficiently large, high-quality datasets. Since FMs like LLMs are trained on enormous amounts of data, applying SBSE to optimize techniques such as prompting or few-shot learning necessitates access to similarly extensive and reliable data. Despite this challenge, an overview of existing works on applying SBSE to LLMs indicates that SBSE has significant potential for optimizing SE tasks utilizing FMs. In summary, this highlights several open problems that need to be addressed and significant opportunities for future research to expand the application of SBSE techniques to a broader range of FMs (e.g., VLMs and MMs) and diverse SE domains. Below, we present possible opportunities and future directions.
4.1 Search-based Code Improvement
@@
4.2 Search-based Testing of FMs-Generated Artifacts
@@
Similar to code generation challenges, the test cases generated from LLMs suffer from several issues, such as erroneous tests, low coverage, and test flakiness [13]. The existing approaches tackle these issues using different LLMs and prompting techniques. Given the well-studied SBST techniques, one potential research direction is to apply SBST to improve tests generated with LLMs. Moreover, in case a large number of tests are generated with LLMs, it is infeasible to run all tests for regression testing in continuous integration/continuous deployment (CI/CD) or DevOps environments. Therefore, another research direction is to apply SBST regression test selection, minimization, and prioritization techniques for LLM-generated tests.
4.3 Search-based Debugging, Repair, and Evolution of FMs-Generated Artifacts

### v2

@@ 4.2.2 Solution Encoding for SBSE Problems
A fundamental step in SBSE is solution encoding for a particular SE problem, to identify the relevant search variables of the problem. This varies from one problem nature to another and requires problem-domain knowledge. FMs with diverse capabilities can support solution encoding for different SE lifecycle phases. For instance, to verify software models with SBSE during the design phase, MMs with both textual and visual capabilities can facilitate solution encoding by analyzing software requirements and design artifacts, often created using graphical modeling languages such as Unified Modeling Language (UML) [126].
4.2.3 FMs for Search Space Definition
Most of the algorithms used in SBSE also require the definition of the search space, i.e., the range of possible values of the search variables. Having a too big search space can hinder the effectiveness of the search; on the other hand, a too small search space could exclude good solutions. FMs can provide a solution in this context. For example, for test case generation, they could identify what areas of the input space of the program should be searched to increase the likelihood of generating tests that cover the whole program. Moreover, they could also identify the minimum granularity of value change that is meaningful, i.e., that can bring an observable change in the program output.
4.2.4 FMs for SBSE Implementation Generation
@@
One advanced application of FMs (LLMs specifically) is code generation. While different approaches [75] have been proposed for LLM-based code generation, no approach is specific to generating implementations of SBSE problems. We forecast that LLMs could be adopted to generate either partial or complete implementations of SBSE problems. In a partial implementation, LLMs can generate code to solve an SBSE problem, which the user then completes manually. Alternatively, LLMs can create complete implementations of SBSE problems that can be further refined. This capability extends across key SBSE components, including solution encoding, operator selection, fitness function design, and search algorithm selection, enabling more efficient and automated problem-solving.
4.2.5 FMs for Testing, Debugging, and Repair of SBSE Implementation
@@
5 FMs for SE Artifacts Generated with SBSE
In this section, we examine the role of FMs in complementing SBSE to enhance SE artifacts. We begin with an overview of the current research landscape, followed by a discussion of open challenges and potential opportunities for future advancements.
5.1 Current Landscape
@@
In recent years, there has been growing research interest in various SE areas in using LLMs for SBSE. Software program refactoring represents another emerging direction where FMs are applied in SBSE. In this regard, Choi et al. [32] utilized LLMs to guide a search-based approach, enhancing refactoring efficiency while reducing complexity. Another key area of interest is software testing, where LLMs are being explored for search-based software testing (SBST). In this case, Turhan [175] introduced evoLLve’M, an approach that utilizes LLMs to enhance JUnit test assertions generated by EvoSuite’s evolutionary process [51, 50]. Similar to this work, Biagiola et al. [17] employed LLMs to improve the readability of unit tests generated by EvoSuite through search techniques.
Search-based techniques have long formed the foundation of REST API testing, gaining significant research attention and leading to the development of robust tools such as EvoMaster [8] and RESTler [14]. Some of these tools have even been applied in industrial contexts, for example, EvoMaster to test healthcare and automotive applications [149, 133]. However, the industrial adoption of these search-based tools was often hindered by their limited capacity for text generation and comprehension [13]. Recently, LLMs, with their advanced natural language processing capabilities, have been increasingly customized and used for REST API testing, giving rise to a new generation of tools such as RESTGPT [83] and LlamaRestTest [84]. Similarly, in the context of evolving applications, search-based regression and differential testing of REST APIs have been initially explored in different domains [58, 73, 153]. LLMs are now being used to perform differential testing on tests originally generated by search-based tools like EvoMaster [74]. This shift represents a major advancement in REST API testing, demonstrating the role of LLMs in enhancing and complementing search-based approaches for modern, dynamic, and evolving applications.
@@
In summary, the review of the literature indicates that (i) LLMs, among various FMs, are increasingly used for SBSE, and (ii) numerous SE phases, including requirements engineering, software design, and debugging, remain largely unexplored.
5.2 Roadmap
Despite SBSE’s long-standing role in solving SE problems, several challenges persist, which we anticipate can be addressed through FMs in the future. Below, we highlight key opportunities and research directions.
5.2.1 FMs for SE Artifact Interpretation
@@
An SE problem in SBSE (e.g., requirements prioritization) is formulated using different solution encoding schemes, such as Binary or Tree-based encoding, depending on the problem’s nature. The final output is generated as an encoded solution, often requiring manual effort for interpretation and transformation into a specific SE artifact. This effort varies based on the SE lifecycle phase. For instance, in the case of requirements prioritization for the next release problem (NRP), the SBSE-generated prioritization sequence needs to be interpreted in textual requirements and utilized for NRP. Similarly, software models generated with SBSE need to be converted to graphical models. In this context, FMs can play a significant role in two directions. First, LLMs with text comprehension and analytical capabilities can be used for SBSE-generated SE artifacts that are required to be interpreted and represented in a textual format. Second, VLMs and MMs can be employed to analyze and transform SBSE-generated SE artifacts into graphical representations for easier interpretation.
5.2.2 FMs for SE Artifact Refinement
@@
For some SE problems, SBSE-generated solutions do not need to be transformed; still, they require manual verification by domain experts. This is particularly evident in scenarios such as test data generation for primitive types and program improvement with genetic programming. In the case of test data generation, test effectiveness using metrics like coverage needs to be analyzed after running tests on the system under test. Similarly, programs generated with genetic programming either need to be compiled and tested, or they need to be manually examined using code inspection and walkthrough techniques. Recent studies have explored the use of LLMs for static analysis, particularly to detect bugs and security vulnerabilities within code [92, 94], as well as to generate static analyzers capable of identifying bug patterns [199]. Beyond these applications, LLMs have significant potential in enhancing artifacts generated with SBSE. For the test data generated using SBSE, LLMs can conduct a static analysis utilizing the source code and test data to provide insights on test effectiveness, such as untested paths or incomplete coverage scenarios. In developing LLM-based static analysis techniques, we foresee the potential for devising novel test effectiveness metrics that complement these techniques. Furthermore, LLM-based static analysis (e.g., type checking and data flow analysis) can be further extended to verify programs generated with SBSE statically, with the potential to significantly reduce false positives through contextual reasoning, as demonstrated in recent studies [92, 94, 199].
5.2.3 FMs for Pareto Analysis and Pruning
@@
In the case of multi- and many-objective SE problems, SBSE generates non-dominated or Pareto-optimal solutions in the objective space. A common practice is to analyze the resulting solutions and select the best one that meets a particular set of requirements. This process needs to be repeated whenever requirements change. In this regard, LLMs can support domain experts in understanding the trade-offs between different objectives and selecting the best solution for a given scenario. Moreover, for non-expert users, VLMs or MMs can construct a visual representation of the Pareto front, illustrating the set of all non-dominated solutions and offering suggestions to assist in selecting the most suitable option. In the case of pruning Pareto optimal solutions, which involves simplifying the solution set for easier decision-making, several challenges highlight the potential role of FMs. One significant challenge is managing high-dimensional problems with more than four conflicting objectives [131]. Another challenge is maintaining solution quality when pruning methods are applied, particularly in complex problems in domains such as robotics [178]. A further challenge lies in identifying the most influential variables from the pruned set and presenting the results in a way that enables intuitive and intelligent decision-making for users [188].
5.2.4 FMs for Validation of the Generated Artifacts
FMs can be used to validate the generated artifacts from various perspectives. For example, in the case of test scenarios generated for cyber-physical systems like ADS (e.g., using genetic algorithms) [90, 173, 102, 134, 103, 25, 55, 26], FMs can help determine whether the generated scenarios are realistic. To this end, some works already aim to assess the realism of generated test scenarios (e.g., [192]). However, these works remain preliminary and are very specific to test scenarios generated using a single testing technique; therefore, more general methods are required. FMs can also be used to improve the quality of generated test scenarios from various perspectives. For instance, they can increase the diversity of scenarios and support the automated removal of redundant ones. To this end, they can help determine the similarity (both syntactic and semantic) between two scenarios, further reducing duplication. Similarly, FMs can also be used to validate other generated SE artifacts, such as models or code produced with genetic programming. These artifacts can be further enhanced by FMs from both functional and non-functional perspectives (e.g., removing redundant model elements or optimizing code for a specific target hardware).
Automated debugging, such as delta debugging, represents a promising opportunity to leverage FMs, particularly LLMs. Although search-based delta debugging has been well-studied in the literature [204], several challenges exist. One challenge is that automatically generated test inputs often reduce overall test coverage due to limited system information [177]. Another persistent challenge is the identification of flaky tests, which is complicated by code dependencies and limitations of traditional methods like abstract syntax trees to detect code changes [164]. Recent research has begun to explore LLM-based approaches for flaky test detection and repair [116, 31], yet the potential of LLMs in the context of delta debugging remains understudied. This opens up an exciting research opportunity to develop techniques based on coding LLMs for context-aware input generation and improved flaky test detection, potentially transforming the automated debugging process.
6 SBSE for FMs
In this section, we explore the potential of SBSE in addressing the limitations of FMs. We begin with an overview of the current state of the art and then discuss the key challenges and opportunities ahead.
6.1 Current Landscape
A growing area of research explores the application of SBSE techniques to optimize FMs-related activities, such as prompt engineering. LLMs, due to their advanced textual capabilities, have been widely applied in prompt engineering [193]. Recent efforts have targeted various aspects of prompt engineering, including optimizing textual prompts using search algorithms such as genetic algorithms and differential evolution to improve prompt quality [196, 35], applying evolutionary techniques for prompt tuning [64], and employing multi-objective optimization methods to refine instructions within prompts [200]. Moreover, research has explored prompt augmentation through search-based data generation, such as using evolutionary techniques to generate instruction data [195]. Another emerging direction is LLM security testing, where search-based techniques are used to generate adversarial prompts to identify vulnerabilities in LLMs [99, 38].
In SE, a notable focus is on estimating software effort, an essential activity in software project management. For this purpose, search techniques have been applied to optimize few-shot learning methods for fine-tuning LLMs to improve their performance in estimating software effort [172]. Another focus area involves applying SBSE techniques to improve image generation quality, which has significant implications for tasks such as automated user interface design. In this direction, Berger et al. [16] presented StableYolo, which applies a multi-objective search to optimize LLM prompts and parameters to generate high-quality and realistic images using tools such as YOLO and Stable Diffusion [16]. Building on this work, GreenStableYolo was introduced, which applies a multi-objective search for efficient LLM inference while maintaining image quality [60]. A recent study introduced search-based prompt refinement techniques to improve LLM-based code generation [169].
6.2 Roadmap
@@
A key limitation noted in the literature is the lack of sufficiently large, high-quality datasets. Since FMs like LLMs are trained on enormous amounts of data, applying SBSE to optimize techniques such as prompting or few-shot learning necessitates access to similarly extensive and reliable data. Despite this challenge, an overview of existing works on applying SBSE to LLMs indicates that SBSE has significant potential for optimizing SE tasks utilizing FMs. Overall, this highlights several open problems that need to be addressed and significant opportunities for future research to expand the application of SBSE techniques to a broader range of FMs (e.g., VLMs and MMs). Below, we present challenges and possible opportunities for further research.
6.2.1 Addressing FM Challenges
A key challenge of FMs is non-determinism, where the same input can produce varying outputs, affecting consistency and reliability. This non-determinism leads to several challenges in applying FMs to real-world use cases. For instance, in continuous integration and continuous delivery (CI/CD) and DevOps, variability in generated configurations may lead to inconsistencies in software releases [153]. Search algorithms, especially those designed to operate under noise and uncertainty [85, 40], offer promising means to manage non-determinism in FMs. One potential research direction is the development of methods to systematically explore input spaces to identify scenarios where model outputs exhibit high variability. This could involve building on traditional techniques such as noise-tolerant fitness functions, sampling strategies, and robust selection mechanisms [136]. Another promising direction for further research is the application of search techniques in identifying optimal configurations for software and hardware resources that minimize variability in model behavior.
6.2.2 Managing Uncertainty in FMs
Uncertainty is an inherent challenge of FMs. Since FMs are trained on vast amounts of data, fine-tuning them also requires a significant amount of data, and the quality of such data (e.g., redundant features or imbalanced data) can be a potential source of uncertainty. For example, VLMs trained/fine-tuned on images containing poor contrast, low brightness, or blurriness can significantly affect model interference. Such uncertainties can significantly affect model interference, rendering these models ineffective for several domain-specific tasks. SBSE can be applied to select an optimal set of features, which will result in effectively managing data-related uncertainty and efficient training/fine-tuning of FMs.
6.2.3 Hallucinations in FMs
A key challenge in FMs is hallucination, where the model generates entirely fabricated or factually incorrect information [139], posing a significant concern in FM applications. As a result, detecting and mitigating hallucinations has become a key focus of ongoing research [29]. In this regard, we foresee significant potential for SBSE to effectively address the challenge of hallucinations in FMs. Specifically, SBSE can complement black-box methods for hallucination detection, as highlighted by the need for such methods [29]. By leveraging a search-based approach, a search algorithm can iteratively generate a diverse population of prompts, collect corresponding responses from FMs, and use a fitness function to continuously assess response variations, enabling more precise hallucination detection.
6.2.4 Fine-tuning FMs
FMs adapted for different SE phases are typically required to be fine-tuned using methods such as few-shot learning and transfer learning. With the continuous evolution of software following SE development models, FMs used in these phases are required to be fine-tuned repeatedly. For instance, whenever software requirements change, the FMs for that particular software need to be fine-tuned to adapt to updated requirements. Given the size of FM parameters and modality, fine-tuning can be highly computationally expensive. In such cases, SBSE can be used to optimize the FMs’ fine-tuning process, including optimizing various activities like selecting hyperparameters and model configurations. This could result in increasing the overall efficiency and streamlining FMs’ adaptation in the SE development lifecycle.
6.2.5 Prompt Optimization
The performance of FMs is heavily influenced by the design of prompts, which are specified in different formats, including textual, visual, or structured formats like domain-specific languages [62]. However, manual prompt engineering is often labor-intensive and prone to errors, which has led to significant interest in automated prompt engineering [93]. Given the vast search space for identifying optimal prompts, prompt optimization has become an active research area [101]. In this regard, several techniques have been explored, including gradient-based methods [207], LLM-assisted optimization [110], example-guided strategies [197], and reinforcement learning [93]. Evolutionary techniques have also been applied, such as genetic algorithms for few-shot textual prompts in LLMs and evolutionary algorithms for one-shot prompts in VLMs [93, 135]. Despite these advancements, a key challenge remains in balancing trade-offs among objectives such as high generation costs, longer execution times, and the risk of overfitting [135]. A promising research direction is casting prompt optimization objectives into multi- or many-objective search problems and analyzing search algorithms’ effectiveness in exploring trade-offs. This will further open pathways to the development of novel search operators and heuristics tailored for automatic prompt optimization.
7 SBSE for SE Artifacts Generated with FMs
In this section, we analyze the role of SBSE in improving SE artifacts generated with FMs. We begin with an overview of the current research landscape, followed by a discussion of open challenges and potential opportunities for future advancements.
7.1 Current Landscape
FMs are increasingly being integrated into various phases of the SE lifecycle, significantly transforming how SE activities are performed [91, 206]. This integration has led to the development of FMs tailored for specialized SE tasks, enabling more efficient and intelligent automation throughout the lifecycle. Prominent examples include LLMs tailored for various coding tasks, such as CodeX [30] designed for code generation and CodeGen [124] for program synthesis. In addition to code generation, LLMs have also been used in software testing to automatically generate test cases, significantly improving the efficiency of testing processes [46]. For debugging, LLMs have demonstrated usefulness for tasks such as fault localization, where they assist in identifying the root cause of software failures [194, 198, 100]. Furthermore, these models have been used to reproduce software bugs by interpreting bug reports and generating the corresponding scenarios using advanced prompting techniques [79].
As LLMs are increasingly used for various SE tasks, there are also several challenges. For the code and test cases generated using LLMs, evaluating their correctness is a key challenge [30, 124]. In addition, since FMs (including LLMs) are trained using open-source code and datasets, they often suffer from a lack of diversity, which is necessary for tasks such as fault localization and program synthesis [194]. Another challenge is that FMs have no direct control over the execution environments of code and tests, thereby limiting their ability to support activities such as bug reproduction [79]. To address these challenges, the SBSE community has begun to apply search optimization techniques to the SE artifacts generated with FMs. Recently, Wang et al. [186] utilized SBSE to optimize LLM-generated code based on the given context. Similarly, Lijzenga et al. [97] applied SBSE to generate code patches to repair programs generated by coding LLMs.
7.2 Roadmap
Despite significant efforts to apply SBSE to enhance SE artifacts generated with FMs, several challenges remain, highlighting the need for further research opportunities.
7.2.1 Search-based Code Improvement
@@
7.2.2 Search-based Testing of FMs-Generated Artifacts
@@
Similar to code generation challenges, the test cases generated from LLMs suffer from several issues, such as erroneous tests, low coverage, and test flakiness [46]. The existing approaches tackle these issues using different LLMs and prompting techniques. Given the well-studied SBST techniques, one potential research direction is to apply SBST to improve tests generated with LLMs. Moreover, in case a large number of tests are generated with LLMs, it is infeasible to run all tests for regression testing in CI/CD or DevOps environments [153]. Therefore, another research direction is to apply SBST regression test selection, minimization, and prioritization techniques for LLM-generated tests. Moreover, during test optimization, search-based testing can improve test diversity by removing redundant test cases. Such research opportunities still exist and require further investigation in the coming years.
7.2.3 Search-based Optimization of Models Suggested by FM
Models generated with FMs, such as various UML diagrams, can potentially be further optimized using search algorithms. For instance, when FMs suggest multiple model variants, search-based techniques could be employed to select the best one based on defined fitness functions. Moreover, constraints defined on models may not be easily tested by FMs alone, and search-based testing techniques can help ensure their correctness. Although methods for such testing exist [76, 21], their integration with FMs remains largely unexplored, and further studies are needed to systematically investigate how these combined approaches can enhance model quality and correctness.
7.2.4 Search-based Optimization of Configurations Suggested by FM
Recent studies have started exploring FMs, specifically LLMs, in identifying software configurations; however, there are several performance challenges [165]. While search algorithms offer a promising approach to optimizing FM-generated software configurations (e.g., test, deployment, or software product line configurations), more work is needed to fully explore their potential in this context. Future research should investigate the effectiveness of different search techniques, the impact of various fitness functions on functional and non-functional characteristics of software, and the integration of FM-generated initial configurations as an initial population for search-based techniques. Such studies could provide deeper insights into how these combined approaches can systematically improve software quality and performance by optimizing software configurations.
7.2.5 Search-based Debugging, Repair, and Evolution of FMs-Generated Artifacts